Cove’s Cakes is a local bakery. Price and cost information follows:

Price per cake $ 14.01
Variable cost per cake
Ingredients 2.23
Direct labor 1.02
Overhead (box, etc.) 0.11
Fixed cost per month $ 4,153.50


Required:
1.
Calculate Cove’s new break-even point under each of the following independent scenarios: (Round your answer to the nearest whole number.)

a. Sales price increases by $2.00 per cake.



b. Fixed costs increase by $500 per month.



c. Variable costs decrease by $0.29 per cake.



d. Sales price decreases by $0.30 per cake.



2. Assume that Cove sold 415 cakes last month. Calculate the company’s degree of operating leverage. (Do not round intermediate calculations. Round your answer to 2 decimal places.)



3. Using the degree of operating leverage calculated in Requirement 2, calculate the change in profit caused by a 13 percent increase in sales revenue. (Round your final answer to 2 decimal places (i.e. .1234 should be entered as 12.34%.))
Answer #1
Req 1.
case a: sales price increased by 2.00 per unit
Cm per unit:
Sales price 16.01
Less: VC per unit (2.23+1.02+0.11) 3.36
Cm per unit: 12.65
Break even units = Fixed cost / CM per unit
4153.50 / 12.65 = 328.34 units
Case b: Fixed cost increase by 500 per month
Seleas price 14.01
Less: VC pe runit 3.36
CM per unit 10.65
Break even units = Fixed cost / CM per unit
(4153.50+500) / 10.65 = 436.95 units
case c: VC decrease by 0.29 per unit
Sales price 14.01
Less: VC per unit (3.36-0.29) 3.07
CM per unit 10.94
Brezk even units = Fixed cost / CM per unit
4153.50 / 10.94 = 379.66 units
Cased: Sales price decrease by 0.30 per unit
Sales price (14.01- 0.30) 13.71
Less: VC per unit 3.36
CM per unit 10.35
\Break even units = Fixed cost / CM per unit
4153.50 /10.35 = 401.30 units
Req 2.
CM per unit = 14.01 - 3.36 = 10.65
Sales units = 415 units
Total contribution (415*10.65) 4419.75
Less: Fixed cost 4153.5
Net income 266.25
Degree of Operting leverage = Total contribution/ Net Income
4419.75 /266.25 =16.6 times
Req 3.
Net Sales increasses by 13%
Multiply: Degree of operating leverage 16.6
Net Income increase by 215.80%